OpenStudy (anonymous):

solve 50 x^5-7x ^4+x ^2 divided by x

OpenStudy (andriod09):

this is the problem, correct? \[\frac {50 \times x^5-7x^2} {x}\]

OpenStudy (anonymous):

if( (50*x^5-7*x^4+x^2)/x)=50*x^4-7*x^3+x. After doing a division, you assume that x!=0(not equal to 0) so you left with 50*x^4-7*x^3+x=x*(50*x^3-7*x^2+1) x*(50*x^3-7*x^2+1)=0 means that either x=0 or (50*x^3-7*x^2+1)=0, since we have assumed that x!=0 just (50*x^3-7*x^2+1)=0 50*x^3-7*x^2+1=0 has only one real solution x= -0.23190047219364127, the others are complex